OpenStudy (anonymous):

can someone help please one angle of a parallelogram measures 30°. what are the measures of the other three angles? a.) 30°,90°,90° b.) 30°,150°,150° c.) 60°,60°,150° d.) cannot be determined

OpenStudy (anonymous):

do you have a diagram

OpenStudy (anonymous):

no there is no diagram sorry

OpenStudy (anonymous):

thats fine

OpenStudy (anonymous):

well a parallelogram has two sets of equal angles so 30+30=60 360-60=300 300/2=150

OpenStudy (anonymous):

30 150 150

OpenStudy (anonymous):

omg thank you so much your a lifesaver thank you

OpenStudy (anonymous):

your welcome!

OpenStudy (anonymous):

thank you again so much

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

im going to close the question now so i can maybe put a different one

OpenStudy (anonymous):

ok im just chillin here being bored